# Exciton self-trapping in twisted hexagonal boron nitride homostructures

### v1

@@
Figure 4: (a) CL spectra normalized by the excitation power for different excitation densities (5kV, 300K) measured on the HPHT-11° sample (b) $IQE$ of the 300 nm emission, deduced from (a), as a function of the excitation power density. The black arrow indicates the transition point between the two methods for varying the excitation density, as indicated in the text. (c) $IQE$ of the 300 nm emission as a function of excitation density for the 5 HPHT samples.
Fig. 4(a) shows the spectra normalized by the excitation power for different excitation densities on the HPHT-11° sample. Fig. 4(b) further shows that the $IQE$ of the 300 nm luminescence, measured on the spectra on (a), decreases by a factor of 30 between 0.1 and 1400 W.cm-2. This decrease reveals a non-linearity in the excitation of the 300 nm emission that appears at extremely low power densities. Fig. 4(c) shows that this saturation effect appears systematically above 10-1 W.cm-2. Note that the twist angle seems to have a minor effect on the saturation threshold. This plot provides a practical guideline for employing CL excitation conditions that avoid non-linear effects in the sample study.
More interestingly, for excitation densities below 0.1 W.cm-2, Fig. 4(c) shows that the $IQE$ of the 300 nm luminescence from the interface can reach surprisingly high values, close to unity. The $IQE$ results are presented in Tab. 2. The higher the twist angle, the higher the $IQE$ of the 300 nm luminescence. The $IQE$ approaching 100% near a 30° twist is a very rare phenomenon at 300 K and may be promising for light emitting devices. This means that (i) the transport of free excitons from the crystal volume to the interface has an efficiency close to 100%, which is not surprising considering the exciton out-of-plane diffusion lengths measured in HPHT samples [61], and (ii) close to 30° twist angle, the mechanism at the origin of the 300 nm luminescence from the interface has an efficiency close to 100%. This means that close to 30°, almost all excitons trapped at the interface recombines with a photon emission within the 300 nm band.
Sample 300 nm $IQE$ (%)
@@
Table 2: $IQE$ of the 300 nm emission measured below 0.1 W.cm-2 on twisted hBN homostructures made of HPHT crystals.
Note that between 10-2 and 10-1 W.cm-2, the average energy of the photons emitted within the broad 300 nm band is slightly red-shifted by about 15 meV when increasing the power. In contrast, a donor-acceptor pair (DAP) emission, which would have been a candidate for such a broad luminescence, shows a clear blue shift with increasing power excitation [62, 63]. This provides a strong evidence that the 300 nm emission of hBN-hBN structures is not due to DAP recombinations.
@@
The characteristics of the 300 nm emission (2 eV energy shift and the 2 eV broadening) supports a self-trapping mechanism of excitons at the interface that we develop in this section. Fig. 5(a) illustrates the difference between simple exciton trapping in the potential well formed by the interface between the two twisted hBN flakes (X2D) and self-trapping of this exciton at the interface (XST). A self-trapped exciton is an exciton that induces a lattice distortion around it, and that is trapped in this distorted area. The configuration diagram drawn in Fig. 5(b) displays the exciton energy as a function of the lattice deformation around it, represented by the configuration coordinate. In this diagram, we define the trapping energy potential of the twisted interface at zero deformation, denoted $E_{T}$, which is estimated in the order of 100 meV according to theoretical calculations [42, 41], as well as an energy barrier to the self-trapping, denoted $E_{FST}$. The possibility of crossing an energy barrier for self-trapping was predicted by Landau [43] and Toyozawa [64] for electrons and by Rashba [65, 66] and Sumi and Toyozawa [67] for excitons. This was demonstrated experimentally in the 60’s and 70’s for self-trapped excitons on alkali halides [68, 69, 70]. The barrier disappears only in the case of self-trapping of an exciton previously trapped on a 0D site [66, 71]. The presence of this energy barrier implies the coexistence of excitons simply trapped at the 2D interface (X2D) and excitons self-trapped at the interface (XST). By adding the free excitons of the 3D bulk hBN (X3D), there are thus three coupled exciton populations during the CL experiments. Note that the energy barrier for self-trapping is necessarily low, since luminescence of XST occurs close to liquid helium temperature (see Appendix E). However, it can limit the formation of XST and reduce the intensity of their luminescence as the temperature is lowered [68, 69, 70]. In this process, radiative recombination of a self-trapped exciton is accompanied by the emission of a set of phonons to restore the crystal lattice to its undistorted form as sketched in Fig. 5(b). This drastically reduces the energy of the emitted photon with respect to the band gap of the interface and explains the large energy shift ($\approx$ 2 eV) between the excitonic bandgap and the luminescence at the interface. Furthermore, the curvature of the ground state at high distortion is such that recombining self-trapped excitons inherently leads to very broad emission. This is illustrated in Fig. 5(b), where the three purple arrows (solid lines) represent photons of different energies resulting from the recombination of self-trapped excitons at the band edge.
@@
Figure 5: (a) Diagram of a trapped exciton at the interface (X2D), in red, and of a self-trapped exciton at the interface (XST), in purple, with a self-induced lattice deformation on which the exciton is trapped. (b) Configuration coordinate diagram representing the three excitons populations in hBN-hBN structures: X3D, X2D and XST. The exciton trapping potential of the interface at zero distortion appears as $E_{T}$ between X3D and X2D. The energy barrier for the formation of the self-trapped exciton is indicated ($E_{FST}$). The long straight arrows represent the photons resulting from the radiative recombination of free excitons (in black), of thermalized self-trapped excitons (purple, solid lines), and of non-thermalized self-trapped excitons or ”hot” self-trapped excitons (purple, dashed line). The small curved arrows represent the phonons required for the lattice to return to the ground state during the recombination of a self-trapped exciton.
@@
Experimentally, emission related to excitons trapped at the interface (X2D) has not been detected in CL. Two possible scenarios can be considered to account for the zero-luminescence intensity: 1) the X2D population is zero, 2) the X2D excitons are dark, i.e. have an infinite radiative lifetime. On one hand, in the first scenario, excitons necessarily accumulate at the interface in the XST state, which, first, requires their self-trapping to be instantaneous, and, second, implies that the detrapping of excitons from the interface into the volume (shown earlier) occurs from the XST states. This seems to contradict the presence of the energy barrier to self-trapping that we emphasized earlier and that limits the formation of XST. Furthermore, detrapping from an XST state to an X3D state requires the emission of a large number of phonons necessary to restore the lattice to its undistorted form, which will be energetically too costly. On the other hand, in the scenario 2, X2D excitons have a negligible probability of radiative recombination (dark excitons). This seems to be consistent with (i) the non-radiative nature of the free surface [61], (ii) the long lifetime of the interface excitons evidenced in Fig. 3, and (iii) the $IQE$ close to 100% observed for XST at high twist angles (Tab. 2), implying that the recombination rate of X2D is very low compared to that of XST. Therefore, we have ruled out scenario 1 and in the following we will only consider scenario 2. Since the bright recombinations are dominant at high angles (see table 2.), we will also neglect the non-radiative recombinations of X2D.
.6 VI - Temperature-dependent investigations.
In order to strengthen the attribution of the 300 nm emission to a self-trapping process, and to characterize the different exciton populations in more details, we have performed a series of experiments as a function of temperature. We first studied the spectral features of the 300 nm luminescence. Fig. 6(a-b) shows the cw-CL spectra recorded on the HPHT-11° sample, at very low power density (5 kV, 27 pA, 60 µm spot diameter) to avoid saturation effects. Since saturation could not be avoided below 100 K even at this very low power density ($\approx$ 0.005 W.cm-2), we limited the measurements to the 100-300 K range.
@@
Non-thermalized (hot) self-trapped excitons are favored at high temperatures, and their recombination, shown in Fig. 6(b), occurs at a higher energy than when they are thermalized to their energy minimum. Theoretically, we therefore expect a slight increase in the emission energy of self-trapped excitons as the temperature increases [72, 73]. Between 200 and 300 K, there is a blue shift of the XST exciton luminescence in the 250-300 nm range, consistent with a contribution from non-thermalized XST that increases at higher temperatures (hot XST in Fig. 6). This broadens the emission towards high energies, as characterized by the increase in the standard deviation of the emission energy (Fig. 6(c)) and the shift of the average energy of the XST emission (Fig. 6(d)). Note that the spectral width of the hBN-hBN 300 nm luminescence still remains around 2 eV at cryogenic temperatures, as shown in the Appendix E. This is in complete contrast to the color center emission that can also occur on some hBN-hBN structures and on single hBN flakes, which exhibit very narrow linewidths at 5K [74]. This clearly indicates that the luminescence from the color centers and from the hBN-hBN interfaces are of different natures.
Figure 6: (a-b) Spectra recorded on the HPHT-11° sample between 100 and 310 K. (c) Standard deviation of the emission energy, (d) average emission energy, and (e) $IQE$ of the self-trapped exciton XST luminescence as a function of temperature. An Arrhenius law is shown in red, from which an activation energy of about 30 meV is extracted.
@@
Fig. 6 (e) shows that the $IQE$ of the XST luminescence drops below 200 K, and is only a few percent at 100 K. The activation of XST emission with temperature is consistent with an energy barrier for self-trapping (named $E_{FST}$ in Fig. 5(b)). We evaluate this energy $E_{FST}$ with an Arrhenius plot fit. We find an $E_{FST}$ value of 31 meV. This value is undoubtedly overestimated due to saturation, which is probably not completely avoided at about 100 K, and is subject to a high uncertainty due to the small temperature range. It is therefore reasonable to estimate the activation of self-trapping in this sample to be around 10 meV. This is the order of magnitude found in the literature for the self-trapping of excitons in other materials such as alkali halides [72, 75, 68, 70, 76]. The various effects revealed by these first temperature-dependent experiments are all consistent with an emission resulting from the recombination of excitons self-trapped at the twisted hBN-hBN interface.
On this basis, we now explain the recombination dynamics of the three exciton populations in hBN-hBN homostructures (X3D, X2D and XST), considering the interactions that bind them, in order to ascribe a physical meaning to the decay times in Fig. 3. The characteristic time constants of the different transitions between the different exciton populations are shown in Fig. 7. X2D accumulate with a rate $C$ at the interface following the efficient capture of the free excitons generated by the primary excitation at the interface. They are then slowly converted, at a detrapping rate $D$, to X3D or self-trapped in the XST states at a rate $ST$.
@@
Figure 7: Scenario of energy levels and transitions between the three exciton populations within the twisted hBN-hBN homostructures. In this scenario, excitons from volume (X3D) are trapped at the interface, where they accumulate in the dark and long living form (X2D). Once self-trapped (XST), they recombine very efficiently with a photon emission, leading to the intense emission at 300 nm.
@@
According to this scenario, the long decay time of X3D and XST measured in TRCL ($\tau_{l}$) after the interruption of the continuous excitation (see Fig. 3) corresponds to the lifetime of the X2D excitons that feed the X3D and XST populations when they detrapped or self-trapped. Considering the detrapping (rate $D$) and self-trapping (rate $ST$) of X2D, $\tau_{l}$ is given by: 1/$\tau_{l}$approx $D$+$ST$ (Details of the calculation are given in the Appendix F.). From this relation, the detrapping rate $D$ can be evaluated from temperature-dependent TRCL measurements of the long decay time of X3D and XST. These decay times are similar for both exciton populations when decreasing the temperature (see Fig. 3 for room temperature), so we present the measurements only for the free excitons luminescence. The results are shown in Fig. 8(a) for HPHT-11° sample. The decays are normalized by the integrated intensity of the X3D band measured by CL spectroscopy under continuous excitation (spectra in Fig. 6(a-b) for sample HPHT-11°). The characteristic time of the long component $\tau_{l}$ is extracted from bi-exponential function fits.
@@
Figure 8: (a) TRCL decays of free exciton luminescence for temperatures in the range 150 - 293K on HPHT-11° sample. Bi-exponential decays are used in the fits to extract the characteristic time of the long component $\tau_{l}$. (b) 1/$\tau_{l}$ as a function of 1/T in the HPHT-11° sample. An Arrhenius law is fitted from which a 110 meV activation energy is deduced.
Fig. 8(b) is a plot of 1/$\tau_{l}$approx $D$+$ST$ as a function of 1/T(K). The data are fitted with an Arrhenius function from which an activation energy of 110 meV is extracted. Given the low activation energy for the formation of self-trapped excitons ($E_{FST}$), previously estimated to be around 10 meV, 1/$\tau_{l}$approx $D$+$ST$ appears to be limited by the thermal activation of the detrapping rate $D$. A comparable activation energy can be extracted from the temperature evolution of $\tau_{l}$ measured on the HPHT-15° sample. We deduce that the exciton trapping potential at the interface is of the order of 100 meV for these samples whose twist angle is of the order of 10-15°. This experimental measurement of a shallow trapping of exciton at the twisted interface is in agreement with previous theoretical calculation [42, 41], and evidence that the 2 eV redshift of the interface luminescence does not come from a giant Moiré trapping, as suggested by previous studies [39, 40].
In summary, we propose a scenario of excitons accumulating at the interface in a non-radiative excitons X2D form and their subsequent self-trapping. This phenomenological description accounts for the observed luminescence decay and is consistent with the luminescence efficiencies close to 100% measured for 30° twist at 300 K for the recombination of self-trapped excitons. It was also concluded that for intermediate twist angles, on the HPHT-11° and HPHT-15° samples, the long decay time between 150 and 300 K is limited by the rate of exciton detrapping from the interface. Thermal activation of detrapping revealed an interface exciton trapping potential of the order of 100 meV for both samples, in agreement with theoretical estimates [42, 41].
.7 VII - Discussion: Self-trapping in sp2 BN
We now discuss the origin of the exciton self-trapping in hBN. First, exciton self-trapping is favored by the ability of the lattice to deform around the exciton, and second, it is facilitated by the ability of the exciton to be localized in the deformed area it has induced. Given the high ionicity of hBN, exciton formation involves the transfer of charge from a boron atom (hole on boron) to a nitrogen atom (electron on nitrogen), as illustrated by theoretical calculations of the intensity of the exciton wave function [77, 78]. This charge transfer induces the formation of a dipole to which the ionic lattice is sensitive. Furthermore, due to the high compactness of the hBN exciton [50], the size of the dipole it forms is comparable to that of dipoles formed by ionic bonds, so the lattice-exciton interaction is relatively strong in hBN. These features are highly beneficial for exciton self-trapping in hBN, yet the latter is not observed in the defect-free crystal. In our study, it appears only after their localization at a highly disoriented interface. On the one hand, at low twist angles, a Moiré superlattice and a resulting periodic superpotential appears and may interact with excitons [18, 19, 20, 21, 22, 23, 24]. On the other hand, at a high degree of twist, a transition from a commensurable to an incommensurable lattice occurs, as studied in quasicrystals, leading to a drastic change in the structure and the physical phenomena it can host. While at low hBN-hBN twist angles, the perturbation of the charge distribution induced by the appearance of the $\pi^{B-B}$ and $\pi^{N-N}$ homo-bonds is weak and confined to certain zones of the Moiré structure; at high angles a very large fraction of the $\pi$-bonds are perturbed and densely distributed all over the interface. As a result, the potential perturbations increase in locality and density with increasing twist angle, which locally boost the ionic character of the lattice, and as a consequence, the exciton-lattice interaction. In addition, the lattice deformability is likely to be locally strongest at high angles. These two elements address the role of the twist angle on exciton self-trapping, but also highlight the need for further theoretical studies to better understand the origin of exciton self-trapping at the twisted interface, by exploring which atomic sites host self-trapped excitons and which lattice distortion are involved.

### v2

@@
Figure 5: (a) CL spectra normalized by the excitation power for different excitation densities (5kV, 300K) measured on the HPHT-11° sample (b) $IQE$ of the 4-eV emission, deduced from (a), as a function of the excitation power density. The black arrow indicates the transition point between the two methods for varying the excitation density, as indicated in the text. (c) $IQE$ of the 4-eV emission as a function of excitation density for the 5 HPHT samples.
Figure 5(a) shows the CL spectra normalized by the excitation power for different excitation densities on the HPHT-11° sample. In Fig. 5(b), it is observed that above 0.1 W/cm2, the efficiency of the 4-eV luminescence decreases significantly, with $IQE$ $\propto$ power-0.38. This corresponds to a sublinear regime in CL intensity with ICL $\propto$ power0.62, suggesting a bimolecular recombination of excitons at high power densities. Figure 5(c) indicates that this saturation effect appears in all HPHT homostructures with little effect of the twist angle on the 0.1 W/cm2 saturation threshold. In contrast, under similar excitation conditions, bulk hBN luminescence only saturates above 50 W/cm2 due to exciton-exciton annihilation (EEA) of free excitons [64] [^1] Experiments are performed at 10 K in ref. [64], and at 300 K in ours. However, since in such samples the diffusivity is temperature independent [63], we do not expect drastic differences in the probability of exciton-exciton collisions and thus in the EEA rate as the temperature is increased.. The much lower threshold in twisted hBN-hBN structures suggests that saturation is rather driven by the EEA of interface excitons. This interpretation is consistent with the much higher exciton density at the interface than in the volume, which results from the efficient capture of excitons at the interface, as shown later in this section.
Interestingly, at low power densities (below 0.1 W/cm2), the $IQE$ of the 4-eV luminescence can reach very high values, as overviewed in Tab. 2. We observe that the $IQE$ increases with the twist angle, reaching nearly 100% at a 30° twist. This suggests that, for 30° twist angles, the vast majority of excitons are trapped at the interface and recombine via photon emission in the 4-eV band. Such a highly efficient emission is rare at 300 K and may be promising for light emitting devices. A unity $IQE$ of the interface luminescence implies that (i) the transport of free excitons from the bulk to the interface is without losses, and (ii) that the exciton recombinations at the interface are fully radiative. Previous results on exciton diffusion in bulk hBN monocrystals suggest that the conversion of 3D excitons into 2D interface/surface excitons is extremely efficient even for dark interfaces (with low twist angles) and for the non-radiative free surface [63]. This study further shows that the recombination of the 2D interface excitons is mostly radiative near the 30° twist angle quasi-crystal limit.
Sample 4-eV $IQE$ (%)
@@
Table 2: $IQE$ of the 4-eV emission measured below 0.1 W/cm2 on twisted hBN homostructures made of HPHT crystals.
Note that between 10-2 and 10-1 W/cm2, the average energy of the photons emitted within the 4-eV band is slightly red-shifted by about 15 meV as the excitation power is increased. In contrast, a hypothetical donor-acceptor pair (DAP) recombination, which would have been a candidate for such a broad emission, should exhibit a blue-shift when increasing the power excitation [66, 67]. Our observations then discard a DAP origin for the 4-eV emission.
As a summary of this section, power-dependent experiments reveal a strong decrease in the efficiency of the 4-eV emission at moderate excitation, probably caused by a bimolecular annihilation process between excitons accumulated at the hBN-hBN interface. At low excitation, in the most favorable cases, i.e., at twist angles close to 30° with HPHT crystals presenting a long-range exciton diffusion, most excitons recombine radiatively at the interface with a 4-eV photon emission.
@@
The characteristics of the 4-eV emission (2-eV energy shift and the 2-eV broadening) supports a self-trapping mechanism of excitons at the interface that we develop in this section.
Figure 6(a) illustrates the difference between simple exciton trapping in the potential well formed by the interface between the two twisted hBN flakes (X2D) and self-trapping of this exciton at the interface (XST). A self-trapped exciton is an exciton that induces a lattice distortion around it, and that is trapped in this distorted area.
@@
Figure 6: (a) Schematics of a trapped exciton at the interface (X2D), in red, and of a self-trapped exciton at the interface (XST), in purple, with a self-induced lattice deformation on which the exciton is trapped. (b) Configuration diagram representing the three excitons populations in hBN-hBN structures: X3D, X2D and XST. The exciton trapping potential of the interface at zero distortion appears as $E_{T}$ between X3D and X2D. The energy barrier for the formation of the self-trapped exciton is indicated ($E_{ST}$). The long straight arrows represent the photons resulting from the radiative recombination of free excitons (in black), of thermalized self-trapped excitons (purple, solid lines), and of non-thermalized self-trapped excitons or ”hot” self-trapped excitons (purple, dashed line). The small curved arrows represent the phonons required for the lattice to return to the ground state during the recombination of a self-trapped exciton.
The configuration diagram drawn in Fig. 6(b) displays the exciton energy as a function of the lattice deformation around it, represented by the configuration coordinate. In this diagram, we define the trapping energy potential of the twisted interface at zero deformation, denoted $E_{T}$. Its value for the twisted interface in hBN is a priori estimated to be in the order of 100 meV following recent theoretical calculations [42, 41].
The energy barrier for the self-trapping phenomenom is denoted $E_{ST}$ in Fig. 6(b). Its existence was predicted by Landau [43] and Toyozawa [68] for electrons and by Rashba [69, 70] and Sumi and Toyozawa [71] for excitons. The presence of an energy barrier for self-trapping was evidenced experimentally in the 60’s and 70’s in alkali halides [72, 73, 74]. The energy barrier disappears only for self-trapping of a localised exciton on a 0D site [70, 75], while here the hBN-hBN interface is a 2D defect.
The presence of the energy barrier $E_{ST}$ implies the coexistence of excitons simply trapped at the 2D interface (X2D) and excitons self-trapped at the interface (XST). Considering the free excitons in the 3D bulk hBN (X3D), there are thus three coupled exciton populations during the CL experiments: X3D, X2D and XST.
Note that, though $E_{ST}\neq 0$, the energy barrier for self-trapping is necessarily low, unless an efficient tunnelling occurs, since luminescence of XST is still observed close to liquid helium temperature (see Appendix E). However, its presence basically limits the formation of XST and reduce the intensity of their luminescence as the temperature is lowered [72, 73, 74].
The general features of self-trapped excitons are well consistent with our observations. Indeed, the radiative recombination of a self-trapped exciton is accompanied by the emission of a set of phonons to restore the crystal lattice to its undistorted form as sketched in Fig. 6(b). This drastically reduces the energy of the emitted photon with respect to the band gap of the interface and explains the large energy shift ($\approx$2 eV) between the excitonic bandgap and the luminescence at the interface. Furthermore, the luminescence of self-trapped excitons is inherently broad for several reasons. First, due to local disorder at the interface (which is likely to be high at large twist angles, see Fig. 1(d)), the variation of the local potential landscape might cause self-trapped excitons to exhibit a variety of energies and distortions. Second, the recombination of self-trapped excitons requires multiple phonon emissions, which further increase the broadening of their luminescence. This is illustrated in Fig. 6(b), where the three purple arrows (solid lines) represent photons of different energies resulting from the recombination of different self-trapped excitons at the band edge.
@@
Experimentally in this study, the emission related to excitons simply trapped at the interface (X2D) has not been detected in CL. Two possible scenarios can be considered to account for an almost zero-luminescence intensity: either 1) the X2D population is empty, or 2) the X2D excitons are dark, i.e. have an infinite radiative lifetime. In the first scenario, excitons would accumulate at the interface in the self-trapped form as a XST state. This requires their self-trapping to be almost instantaneous, which contradicts the presence of an energy barrier to self-trapping limiting the formation of XST. Scenario 1) also implies that the detrapping of excitons from the interface into the volume occurs from the XST states, which would be energetically too costly. The first scenario clearly appears inconsistent with the significant release of free excitons in the bulk hBN volume from the interface observed by TRCL. Therefore, we have ruled out scenario 1) and in the following we will only consider scenario 2).
In this scenario, X2D excitons have a negligible probability of radiative recombination (dark excitons). This is in good qualitative agreement with (i) recent investigations [63] which show that the free surface of hBN single crystals acts as a non radiative trap for excitons, (ii) the long lasting release of the interface excitons in the volume, and (iii) the $IQE$ close to 100% observed for XST at high twist angles (Tab. 2), implying that the recombination rate of X2D is very low compared to that of XST. Since the IQE reaches almost unity at high angles (see table 2.), the localised self-trapped excitons then recombine very fast and radiatively, with an almost instantaneous radiative recombination rate compared to the other processes that have been considered. For this reason, we will further neglect the eventual non-radiative recombinations of X2D. Figure 8 further summarizes these conclusions.
.6 VI - Temperature-dependent experiments
To strengthen the attribution of the 4-eV emission to a self-trapping process, and to characterize the different exciton populations in more detail, we have performed a series of experiments as a function of temperature.
The intermediate twist angle samples HPHT-11° and HPHT-15° are studied at a very low power density (5 kV, 27 pA, 60 µm spot diameter; $\approx$ 0.005 W/cm2), which prevents saturation effects down to 100 K. Since saturation effects appears to be stronger at low temperature and could not be be avoided below 100 K, we limited the data analysis to the 100-300 K range. We first studied the influence of temperature on the spectral features of the 4-eV luminescence with Fig. 7(a-b).
@@
Non-thermalized (hot) self-trapped excitons are known to be favored at high temperatures, where their recombination, shown in Fig. 6(b), occurs at a higher energy than when thermalized. We therefore expect an increase in the emission energy of self-trapped excitons as the temperature increases [76, 77]. Between 200 and 300 K, Fig. 7(a-b) indeed depicts an increase of the XST exciton luminescence in the 4-5 eV region with respect to the 3-4 eV region, consistent with a contribution from non-thermalized XST. The broadening of the 4-eV emission on the high-energy side, indicated by black arrows in Fig. 7(a-b), is quantitatively characterized by the standard deviation of the emission energy (Fig. 7(c)) and by the average energy of the XST emission (Fig. 7(d)). Both slightly increase at high temperature. Note that the spectral width of the hBN-hBN 4-eV luminescence still remains around 2 eV at cryogenic temperatures, as shown in the Appendix E. This rules out the possibility of color center emission which would exhibit narrow linewidths at 5K [78, 79].
@@
Figure 7(e) shows that the $IQE$ of the XST luminescence drops below 200 K, and is only a few percent at 100 K. The activation of XST emission with temperature is consistent with the presence of an energy barrier for self-trapping as discussed in the previous section (named $E_{ST}$ in Fig. 6(b)). With an Arrhenius plot fit, we find $E_{ST}$ values of 32 and 41 meV for HPHT-11° and HPHT-15° samples respectively. These values are probably overestimated due to the saturation effects evidenced in Section IV, which are not completely avoided near 100 K. They are also subject to a high uncertainty due to the small temperature range investigated. It is therefore reasonable to estimate the activation energy of the self-trapping to be a few tens of meV in these samples. This corresponds well to the order of magnitude found in the literature for the self-trapping of excitons in other materials such as alkali halides [76, 80, 72, 74, 81].
Figure 7: CL spectra recorded on the HPHT-11° (a) and HPHT-15° (b) samples between 100 and 293 K. (c) Standard deviation of the 4-eV emission energy, (d) average emission energy, and (e) $IQE$ of the self-trapped exciton XST luminescence as a function of temperature, extracted from the spectra in (a-b). Arrhenius laws are shown in full lines in (e), from which activation energies are extracted.
As a summary, the various effects revealed by the temperature-dependent experiments appear to be well consistent with a 4-eV emission band resulting from the recombination of excitons that are self-trapped at the twisted hBN-hBN interface.
.7 VII - Phenomenological model for exciton recombination dynamics
@@
Figure 8: Energy levels and transitions between the three exciton populations within the twisted hBN-hBN homostructures. In this scenario, excitons from volume (X3D) are trapped at the interface, where they accumulate in the dark and long living form (X2D). Once self-trapped (XST), they recombine instantaneously with a photon emission, leading to the intense light emission around 4 eV.
We now explain the recombination dynamics of the three exciton populations in hBN-hBN homostructures (X3D, X2D and XST), considering the interactions that bind them, in order to ascribe a physical meaning to the decay times measured in TRCL. Different transitions between these exciton populations are shown in Fig. 8 as a summary of section V conclusions. X2D accumulate at the interface after their capture at a rate $C$ from free excitons of the hBN bulk crystals generated by the primary excitation. They are then either released to the volume as X3D, with a detrapping rate $D$, or self-trapped in the XST states with a rate $ST$.
@@
According to this scenario, the long decay time of X3D and XST measured in TRCL ($\tau_{l}$) after the interruption of the continuous excitation (see Fig. 4) corresponds to the lifetime of the X2D excitons decaying either by detrapping or self-trapping. Considering the associated rates, $\tau_{l}$ is given by: 1/$\tau_{l}\approx D+ST$ (details given in the Appendix F). In the following we discuss the temperature-dependence of the long decay time common to X3D and XST. Since these decay times are found equal for X3D and XST populations when decreasing the temperature, we present only the TRCL measurements of X3D. Fig. 9(a) shows the decays for two samples with intermediate twist angles: HPHT-11° and HPHT-15°. The decays are normalized by the absolute integrated intensity of the X3D band obtained by independant CL measurements under continuous excitation (spectra in Fig. 7(a-b)). The characteristic time of the long component $\tau_{l}$ is extracted from bi-exponential function fits.
@@
Figure 9: (a) TRCL decays of the free exciton luminescence for temperatures in the range 150 - 293 K on HPHT-11° and HPHT-15 samples. Bi-exponential decay fits are used to extract the characteristic time of the long component $\tau_{l}$. (b) 1/$\tau_{l}$ as a function of 1/T in HPHT-11° and HPHT-15° samples. Arrhenius laws appear as solid lines.
Figure 9(b) shows a plot of 1/$\tau_{l}$approx $D$+$ST$ as a function of 1/T(K). The data are fitted with Arrhenius functions, yielding activation energies of 100 and 110 meV for HPHT-11° and HPHT-15° samples, respectively. Given the low activation energy estimated for the formation of self-trapped excitons ($E_{ST}\simeq$ 10 meV), 1/$\tau_{l}$approx $D$+$ST$ is likely limited instead by the thermal activation of the detrapping rate $D$, with an activation energy corresponding to $E_{T}$. The self-trapping of an exciton requires the assistance of many phonons for the distortion to occur. Despite a relatively low activation energy, the self-trapping may be extremely slow at room temperature, in good agreement with next results in this section. It indicates that $D\gg ST$ between 150 and 300 K, consistently with the relatively low efficiency of the 4-eV luminescence observed in these samples at 300 K (see Tab. 2). We conclude that the extracted activation energies correspond to the depth of the potential well experienced by bulk excitons at the twisted hBN interface, ET, which is found to be around 100 meV for twist angles between 10 and 15°. This value is in good agreement with the first theoretical estimates of ET[42, 41].
Coming back to the high twist angles, we remind that the efficiency of the 4-eV emission is close to 100% near 30° twist angles at 300 K (see Tab. 2). This requires an almost full conversion of $X_{2D}$ into $X_{ST}$ states via self-trapping. To achieve this, the detrapping rate is necessarily negligible compared to the self-trapping rate $D\ll ST$ , and the lifetime of the interface excitons is then limited by the self-trapping 1/$\tau_{l}\approx ST$. The situation is opposite to the 10-15° twisted hBN crystals, where 1/$\tau_{l}\approx D$. This comparaison highlights that the interpretation of the long decay time observed in TRCL is cautious depending on twist angles.
Figure 10: (a) TRCL decays of the free exciton luminescence (in black) as compared to the self-trapped exciton luminescence (in purple) for an intermediate twist angle sample HPHT-15° and a high twist angle sample HPHT-27°. Experiments done at 300 K.
We further analyse and compare the TRCL results between 15° and 27° twists at 300 K in Fig. 10. At a 15° twist angle, we then attribute the long decay time to the detrapping rate $D$, found equal to 2.5*106s-1 at 300 K while the self-trapping rate $ST$ is much lower but unknown. At 27° twist angle, we access the self-trapping rate $ST=$2.5*105 s-1, while the detrapping rate $D$ is much lower but unknown. These limit cases are very interesting compared to intermediate twist angles where both should contribute to the long decay time with 1/$\tau_{l}$approx $D$+$ST$.
The self-trapping process is found to be particularly slow around 30° twist (3.2 µs on HPHT-27°, 4 µs on HPHT-24°) though giving high efficiency of the 4-eV emission. Around 15° twist, the dominant detrapping is about 10 times faster (0.40 µs on HPHT-15°, and 0.46 µs on HPHT-11°) with a typical efficiency decrease of the 4 eV emission by a factor of 2 (see Tab. 2). This suggests that the detrapping of excitons from the interface is decreasing at high angle, giving enough time for exciton trapped at the interface to slowly deform the crystal locally around it for self-trapping. This means that the depth of the potential well formed by the twisted hBN interface, ET, found equal to 100 meV at 15°, increases near 30° so that detrapping becomes negligible in front of self-trapping. Theoretical calculations of ET for twist angles between 15 and 30° would be welcome to confirm that point.
In summary, we interpret our experimental data with excitons accumulating at the interface in a non-radiative and long-lived form (X2D), followed by their subsequent conversion into $X_{3D}$ and $X_{ST}$ via their detrapping or self-trapping. Our results suggest that the depth of the interface trapping potential for free excitons in hBN increases with the twist angle. For low angles the detrapping is believed to be dominant and limits the efficiency of the interface luminescence, while close to 30°, the detrapping is negligible and the efficiency of the interface emission reaches almost 100%. Our phenomenological model provides a good description of full the experimental data set, including the long-lasting luminescence and the high $IQE$.
.8 VIII - Discussion: Self-trapping in sp2 BN
We finally discuss the origin of the exciton self-trapping in hBN. Exciton self-trapping is favored by the ability of the lattice to deform around the exciton. Given the high ionicity of the covalent B-N bonds, a strong exciton-lattice interaction is expected in hBN. In the excitons of bulk hBN, the electron and the hole distance are distant from $\AA$ [50], which makes the hBN excitons particularly small. Still an exciton dipole of this size is probably too large to interact with the atomic bonds in bulk hBN and to give rise to self-trapped excitons.
The exciton behaviour at hBN/hBN interface still needs further investigations, but we can already draw general lines. When an exciton is spatially confined, it is a general trend that its binding energy increases and its size decreases. Theoretical calculations have been reported for the free surface of hBN [82] and for the BN monolayer. In the latter, the exciton size is calculated at $\AA$ [83]. A similar trend is expected for 2D excitons confined at the hBN/hBN interface. We then suggest that the exciton size decreases when confined at the interface with respect to the bulk, sufficiently to match the order of the lattice parameter (2.5 $\AA$ in plane, 3.3 $\AA$ out of plane), then promoting the exciton-lattice coupling at the interface by dipole-dipole interactions.
For small twist angles of a few degrees, the interfaces are known to be energetically stabilized by Moiré formation and are probably weakly deformable after that. Indeed, at low angles, intense atomic reconstruction occurs and mechanically stabilizes the structure by expanding the region with stable stacking configurations and condensing all the strain and the unstable stacking configurations into 1D lines [33, 34]. On the contrary, near 30° of twist where self-trapping is better observed, a transition from a commensurable to an incommensurable interface occurs, as studied in quasicrystals, leading to a drastic change in the structure and the physical phenomena it can host. While at low angles the perturbation of the charge distribution induced by the appearance of the $\pi^{B-B}$ and $\pi^{N-N}$ homo-bonds is weak and confined to certain zones of the Moiré structure, at high angles a very large fraction of the $\pi$ bonds are perturbed and densely distributed throughout the interface. As a consequence, the potential perturbations increase in locality and density with increasing twist angle, which could promote a higher barrier for exciton detrapping and favor the slow self-trapping process.
These elements address the role of the twist angle on exciton interactions with the interface, but also highlight the need for theoretical studies to better understand the origin of exciton self-trapping by exploring (i) what kind of hBN lattice distortions are involving self-trapped excitons and (ii) how the potential well seen by hBN excitons does evolve between 5 and 30° twisted interface.